OpenStudy (anonymous):

Need some help please. x = Pi Convert the Cartesian equation to polar equation and sketch the graph. I know that x = r cos (theta) and y = r sin(theta)

OpenStudy (anonymous):

\[r \cos(\theta)+0 \times r \sin(\theta) = \pi\] what \[\theta for \cos = \pi?\] \[\arccos(\cos(\theta) = \arccos(\pi) \]\[\cos^-1(\cos(\theta) = \cos^\] x = -1 y = 0 Is this correct?

OpenStudy (amistre64):

x=pi is a vertical line thru x=pi, right?

OpenStudy (anonymous):

yeah

OpenStudy (anonymous):

how do i turn to polar equation?

OpenStudy (amistre64):

hmm, i vaguely recall a rule for it. let me see if I can relocate it

OpenStudy (anonymous):

okay. thanks!

OpenStudy (amistre64):

http://tutorial.math.lamar.edu/Classes/CalcII/PolarCoordinates.aspx towards the bottom: r cos(t) = pi ; i beleive

OpenStudy (amistre64):

makes sense since x = r cos(t)

OpenStudy (anonymous):

yes. that does make sense. so the y = 0?

OpenStudy (amistre64):

the cartesian form of x=4 has no y part to worry about does it?

OpenStudy (amistre64):

well, x = pi ... has no y parts to consider

OpenStudy (anonymous):

so the graph is just the picture of a vertical line at \[x=\pi\]?

OpenStudy (amistre64):

yes

OpenStudy (amistre64):

http://www.wolframalpha.com/input/?i=x%3Dpi

OpenStudy (anonymous):

So what is the polar equation? \[x=r \cos(\pi)\] ?

OpenStudy (amistre64):

there are no xs in polar equations ...

OpenStudy (anonymous):

Can I ask you another one? \[4x ^{2}+9y^{2}=16\] Same thing.

OpenStudy (amistre64):

x = pi is a cartesian equation rcos(t) = pi is its equivalent polar equation

OpenStudy (amistre64):

hmm, an ellipse in polar form eh

OpenStudy (anonymous):

thinking take sqrt of everything?

OpenStudy (anonymous):

need to turn to polar again...

OpenStudy (amistre64):

there is something more to do with eccentricity and stuff for those http://mysite.du.edu/~jcalvert/math/ellipse.htm has a nice little rundown of it

OpenStudy (anonymous):

okay. thank you sir.

OpenStudy (amistre64):

wish i was more proficient with these, but good luck :)

OpenStudy (anonymous):

prob gonna repost it. thanks!